# A comprehensive framework for statistical testing of brain dynamics

### v1

@@ Abstract
We introduce a comprehensive statistical framework for analysing brain dynamics and testing their associations with behavioural, physiological and other non-imaging variables. Based on a generalisation of the Hidden Markov Model (HMM) — the Gaussian-Linear HMM — our open-source Python package supports multiple experimental paradigms, including task-based and resting-state studies, and addresses a wide range of questions in neuroscience and related scientific fields. Inference is carried out using permutation-based methods and structured Monte Carlo resampling, and the framework can easily handle confounding variables, multiple testing corrections, and hierarchical relationships within the data. The package includes tools for intuitive visualisation of statistical results, along with comprehensive documentation and step-by-step tutorials to make it accessible for users of varying expertise. Altogether, it provides a broadly applicable, end-to-end pipeline for analysis and statistical testing of functional neural data and its dynamics.
@@ Introduction
A fundamental aspect of neuroscience is the necessity to find, assess, and characterise associations between brain activity and behaviour[1, 2]. Understanding these brain-behaviour relationships is crucial for advancing our knowledge of cognitive functions, mental health, and neurological disorders[3, 4, 5]. To quantify such associations, researchers typically rely on either prediction techniques or statistical testing. While prediction methods focus on out-of-sample accuracy, explanatory approaches rely on statistical testing to determine whether specific patterns of interest, such as the relationship between a measurement obtained from the brain and a behavioural trait, are supported by the data[6]. Here, we present easy-to-use and well-documented routines for statistical testing between brain dynamics and different forms of behaviour. These tools can readily be applied in other domains.
The recently introduced Gaussian-Linear Hidden Markov Model (GLHMM) Python package was developed to characterise brain dynamics across multiple data modalities through the identification of latent brain states[7]. This package includes a versatile statistical testing framework that supports a wide range of experimental designs commonly used in neuroscience research, including the resting state, often available in public datasets. With an emphasis on statistical testing, this paper provides a detailed overview of the entire analysis protocol, from model fitting to result presentation.
We define four families of statistical tests that address a wide range of relevant questions in neuroscience. These are: (1) across-subjects tests, which assess associations between individual traits and brain activity across-subjects; (2) across-trials tests, which compare brain activity across trials under different experimental conditions; (3) across-session within-subject tests, which evaluate long-term changes in brain dynamics across multiple sessions for one subject; and (4) across-state-visits tests, which examine associations between brain time series and one or more simultaneously measured variables, such as physiological measurements.
Unlike existing frameworks for statistical inference that primarily target time-averaged or non-temporal data, the present protocol has a strong focus on the temporal dimension of brain activity (i.e. on brain dynamics), even though it can also handle more conventional tests. While these tests are presented with a focus in neuroscience, they can readily be adapted to other fields such as economics, ecology, and other fields. This makes the toolbox suitable for any domain that involves statistical testing of relationships between dynamic system properties (e.g., sequential or temporal data) and a set of external variables.
@@
As represented in Figure 1, the four types of tests presented in this protocol are: across-subjects, across-trials-within-session, across-session-within-subject, and across-state-visits. We next succinctly describe the four tests, and further details can be found in the SI.
Figure 1: Illustration of the four statistical tests. (A) the across-subjects test compares behavioural measurements (or traits) across multiple individuals to test subject trait differences. (B) The across-trials test assesses differences in brain responses across experimental conditions, such as two types of stimuli. (C) The across-session-within-subject test assesses changes in brain responses over experimental sessions, given an experimental paradigm such as the one used in B. (D) The across-state-visits test assesses relationships between state time courses and concurrently recorded signals, where each state may correspond to the activation a specific brain network.
@@
The across-sessions within-subjects test provides a new way to assess whether the brain-behaviour relationship under study changes over slower time scales, i.e. over the course of multiple sessions. This approach can be used in longitudinal studies where a subject is scanned repeatedly while performing a task involving one or more contrasts (e.g. stimuli or subject decisions). Unlike traditional approaches that shuffle trial data, this method operates at the level of regression coefficients. For each session, a regression model is fitted to estimate session-specific beta coefficients, which capture the relationship between brain activity (here, state time courses) and the experimental condition. To test for significant changes across sessions, the method generates a (null) distribution by randomly permuting these beta coefficients across sessions (rather than permuting the data). This approach accounts for differences in session length, variations in condition proportions, and the lack of direct alignment between trials across sessions (see S1 for more details). Similar to the across-trials test, this can be performed at each time point to produce a time-resolved statistical analysis.
The across-state-visits test, also novel in this context, evaluates whether the state time courses (represented by the Viterbi path, a discretised version of the state activation probabilities $\gamma_{tk}$) are associated with concurrently recorded physiological or behavioural signals over time, such as pupil size, heart rate, or skin conductance. For instance, one state might correspond to an increase in pupil size, while another might correspond to a decrease. The Viterbi path thus serves as the contrast, enabling the comparison of differences in the second set of signals. Standard permutation methods are however not suitable for this test, as shuffling time points would disrupt the temporal structure of the data. Instead, using a Monte Carlo approach, we generate surrogate Viterbi paths that preserve the original transition timings but randomly reassign which states are visited at each transition in a structured manner. This is done such that it maintains the statistical properties of the original data while breaking the observed association between states and the external signal, with further elaboration on this test available in S2.
@@
Load and prepare data: Step 4 - Set-up and train a HMM (optional)
@@
As stated, this protocol supports the following types of tests: across-subject, across-trials, across-sessions within-subject, and across-state-visits.
@@
For across-subjects, across-session-within-subject, and across-trials-within-session tests, the protocol provides multivariate regression tests and univariate tests as well as Canonical Correlation Analysis (CCA). See Part 2 in Figure 2.
@@
For the Across-subjects, Across-session-within-subject, and Across-trials-within-session tests, the protocol includes the Non-Parametric Combination (NPC) algorithm to combine multiple p-values into fewer p-values with increased statistical power [24, 25]. Specifically, instead of getting a p-value for each pair of variables—that is a ($p\times q$) matrix of p-values—the NPC algorithm condenses these into either one p-value per row ($1\times p$ p-values), or one p-value per column ($1\times q$ p-values), or a single p-value for the entire test. In our implementation, we use Fisher’s method as the combining function, which efficiently aggregates the p-values while maintaining sensitivity to small values. This approach differs from CCA, which also produces a single p-value but does so by testing the strength of a multivariate relationship between two variable sets. In addition to statistical inference, CCA provides a latent representation of the data in the form of canonical variables—linear combinations that maximise correlation between the sets. NPC, by contrast, aggregates test results across multiple comparisons. While it offers flexibility and interpretability, it does not yield a latent representation or model the joint multivariate structure directly.
@@
This work presents a host of tools for statistical testing of brain dynamics in neuroimaging and neurophysiology. Unlike mainstream toolboxes like FSL, SPM, FreeSurfer, and MNE, whose functionality for statistical testing primarily focus on applying General Linear Models (GLMs) for traditional task-based designs, this protocol supports a wider range of experimental designs, including resting-state paradigms. While solutions for testing in resting-state paradigms exist in other toolboxes, such as functional connectivity approaches (FSL, SPM), surface-based connectivity (FreeSurfer), and spectral analyses (MNE), these approaches typically rely on time-averaged statistics (“static” models) and do not account for temporal variability. Our protocol directly addresses this limitation by enabling dynamic analyses that capture time-varying brain activity and its relationship with behavioural or physiological measures.
Across-subjects and Across-trials analyses are available in FSL, SPM, FreeSurfer, and MNE to perform group-level comparisons and the testing of experimental conditions. However, they do not support across-sessions analyses for studies tracking changes over time, such as longitudinal studies. This protocol uniquely supports Across-sessions testing, as well as across-state-visits testing. The latter examines relationships between dynamic brain states and concurrently measured signals, such as physiological data, using methods like OSR and OSA.
All of the aforementioned toolboxes support both univariate and multivariate tests. However, while FSL, SPM, and FreeSurfer can extend their functionality to include CCA through external tools like Permutation Analysis of Linear Models (PALM), this requires additional setup. MNE lacks this capability entirely. In contrast, CCA is fully integrated into this toolbox, making it easier to use without additional setup.
Our framework also provides hierarchical permutation testing to account for family relationships between subjects. This functionality is fully integrated into the code using PALM, with no requirement for external input or additional installation. MNE does not directly support hierarchical testing.
Regarding multiple testing correction and cluster statistics, all tools support standard approaches, including FWER and FDR. However, this protocol incorporates these corrections directly into its workflow, to streamline their application. Table 2 summarises the main statistical testing capabilities of this protocol in comparison with FSL, SPM, FreeSurfer, and MNE.
Category Feature FSL SPM FreeSurfer MNE-Python
Type of Tests across-subjects Yes Yes Yes Yes
across-trials Yes Yes Yes Yes
@@
No No No No
across-state-visits No No No No
Methods Supported Multivariate Yes (via randomize tool) Yes (e.g., MANOVA) Yes Yes
Univariate Yes Yes Yes Yes
@@
No
one-state-vs-the-rest No No No No
one-state-vs-another-state No No No No
@@
No
Multiple testing correction & Cluster statistics
Family-Wise Error Rate
Yes Yes Yes Yes
False Discovery Rate
Yes Yes Yes Yes
Cluster-based statistics Yes Yes Yes Yes
Table 2: Comparison of the statistical testing capabilities of GLHMM with FSL, SPM, FreeSurfer, and MNE. Each row represents a specific feature or analysis type, with ’Yes’ indicating that the toolbox supports the functionality and ’No’ indicating that it does not support the functionality.
@@
This protocol outlines pipelines for a comprehensive set of statistical tests, applicable to a broad range of scientific questions in neuroscience. We demonstrate these tests using publicly available data, as summarised below. In Protocol 1 (across-subjects), we analyse resting-state brain activity from 1,001 Human Connectome Project (HCP) participants across four sessions to examine its relationship to 15 traits related to cognitive performance, refer to Table S1. In Procedure 2 (across-trials), we study MEG data from a single person who participated in 15 sessions. During each session, the person watched both animate and inanimate objects while their brain activity in the occipital lobe was recorded. This analysis assesses differences in the brain responses when the person looks at animate objects compared to inanimate ones. In Procedure 3 (across-sessions within-subject), we use the same dataset as in Procedure 2, but this time focus on changes over multiple sessions. This analysis shows whether the person exhibits changes in stimulus processing over time (i.e. across sessions) due for example, to learning, or if their brain representations remain stable. In Procedure 4 (across-state-visits), we analyse MEG data from 10 subjects scanned at rest in a dark room. During the scans, pupil size and brain activity were measured concurrently. Nine subjects completed two sessions, and one completed a single session. This analysis explores how changes in brain states, like the default mode network, relate to variations in pupil size[29, 30]. All data (except HCP) needed to reproduce the results of these workflows are hosted on Zenodo (DOI: 10.5281/zenodo.15213970), while the code is available at GitHub.
@@
  • GLHMM Python package: Install the GLHMM Python package and its dependencies using pip. The package is available for download at https://github.com/vidaurre/glhmm. This protocol is based on the latest release of GLHMM (version 1.0.0, released on April 7, 2025).
@@ 1. Install and set up Python environment Timing: 1–5 min
    conda create --name glhmm_env python=3.9
@@
    This creates an environment with Python version 3.9 or newer. Once the environment is activated, clone the GitHub repository and install the GLHMM package using:
    pip install git+https://github.com/vidaurre/glhmm
    The GitHub repository includes all the code required to run the GLHMM package, along with helpful examples and resources. It also contains a requirements.txt file listing additional tools and packages. Run the following command to install these:
    pip install -r requirements.txt
    Datasets and Jupyter notebook tutorials for each of the four procedures outlined in this paper can be downloaded from Zenodo. The data is available under the following DOI: ’10.5281/zenodo.15213970’. You can manually download the data from this DOI and save it in the same folder as the procedures. Alternatively, the provided Jupyter notebooks are designed to automatically download the data if it is not already present in the expected directories.
@@
    Behavioural data for 15 cognitive traits related to fluid intelligence is stored in data_cognitive_traits_HCP.npy and loaded into R_data, referred to as matrix $R$. Look at Table S1 for the full list of traits. Additional confounds, such as sex, age, and fMRI head motion, are included in confounds_HCP.npy.
@@
Procedure 3: Across-session-within-subject testing
Steps 1–5 follow the same process as in Procedure 2, as this protocol uses the same data and preprocessing setup. For details on data loading, preparation, and HMM training, refer to Procedure 2. This protocol starts at Step 6 and focuses on the statistical analysis for the across-sessions within-subject test.
@@
All data used in this protocol are freely accessible on Zenodo (DOI: 10.5281/zenodo.15213970). The Zenodo repository includes references to the GitHub repository for associated code.
@@
All code is available on GitHub as Python notebooks and can be run for free in the cloud using Google Colab. To ensure reproducibility, the GitHub repository is also archived on Zenodo (DOI: 10.5281/zenodo.15213970). This setup supports versioning to track future updates, including new code and tutorials.

### v2

@@ Abstract
We introduce a comprehensive statistical framework for analysing brain dynamics and testing their associations with behavioural, physiological and other non-imaging variables. Based on a generalisation of the Hidden Markov Model (HMM) — the Gaussian-Linear HMM — our open-source Python package supports multiple experimental paradigms, including task-based and resting-state studies, and addresses a wide range of questions in neuroscience and related scientific fields. The toolbox is available as both a Python library and a graphical interface so it can be used by researchers with or without programming experience. Statistical inference is performed using permutation-based methods and structured Monte Carlo resampling, and the framework can easily handle confounding variables, multiple testing corrections, and hierarchical relationships within the data, among other features. The package includes tools for intuitive visualisation of statistical results, along with comprehensive documentation and step-by-step tutorials. Altogether, it provides a broadly applicable, end-to-end pipeline for analysis of functional neural data and its dynamics.
@@ Introduction
A fundamental aspect of neuroscience is the necessity to find, assess, and characterise associations between brain activity and behaviour[1, 2]. Understanding these brain-behaviour relationships is crucial for advancing our knowledge of cognitive functions, mental health, and neurological disorders[3, 4, 5]. To quantify such associations, researchers typically rely on either prediction techniques or statistical testing. While prediction methods focus on out-of-sample accuracy, explanatory approaches rely on statistical testing to determine whether specific patterns of interest, such as the relationship between a measurement obtained from the brain and a behavioural trait, are supported by the data[6].
Here, we present a protocol and associated Python package that provides easy-to-use routines for statistical testing of the relation between brain dynamics and behavioural or physiological variables. These tools are well-documented and easily generalisable to other types of data besides neuroscience. The protocol builds on the Gaussian-Linear Hidden Markov Model (GLHMM) Python package[7]], which implements the GLHMM; this generalizes multiple types of HMMs into a single framework that allows for existing and new models. Based on a HMM-based characterisation of the data, the presented statistical framework supports a wide range of experimental designs commonly used in neuroscience, including the resting state. To support broader accessibility, it also includes a graphical user interface (GUI) that enables users to run analyses without the need of writing code. In this paper, we describe the full protocol, from model fitting to result presentation, with implementation details and worked examples across modalities and experimental designs.
We define four families of statistical tests that address a wide range of relevant scientific questions. These are: (1) across-subjects tests, which assess associations between individual traits and brain activity across-subjects; (2) across-trials tests, which compare brain activity across trials under different experimental conditions; (3) across-sessions-within-subject tests, which evaluate long-term changes in brain dynamics across multiple sessions for one subject; and (4) across-state-visits tests, which examine associations between brain time series and one or more simultaneously measured variables, such as physiological measurements.
Unlike existing frameworks for statistical inference that primarily target time-averaged or non-temporal data, this protocol has a strong focus on the temporal dimension of brain activity (i.e. on brain dynamics), even though it can also handle more conventional tests. While these tests are presented with a focus in neuroscience, they can readily be adapted to other fields such as economics, ecology. This makes the toolbox suitable for any domain that involves statistical testing of relationships between dynamic system properties (e.g., sequential or temporal data) and a set of external variables.
@@
As represented in Figure 1, the four types of tests presented in this protocol are: across-subjects, across-trials-within-session, across-sessions-within-subject, and across-state-visits. We next succinctly describe the four tests, and further details can be found in the SI.
Figure 1: Illustration of the four statistical tests. (A) the across-subjects test compares behavioural measurements (or traits) across multiple individuals to test subject trait differences. (B) The across-trials test assesses differences in brain responses across experimental conditions, such as two types of stimuli. (C) The across-sessions-within-subject test assesses changes in brain responses over experimental sessions, given an experimental paradigm such as the one used in B. (D) The across-state-visits test assesses relationships between state time courses and concurrently recorded signals, where each state may correspond to the activation a specific brain network.
@@
The across-sessions-within-subjects test provides a new way to assess whether the brain-behaviour relationship under study changes over slower time scales, i.e. over the course of multiple sessions. This approach can be used in longitudinal studies where a subject is scanned repeatedly while performing a task involving one or more contrasts (e.g. stimuli or subject decisions). Unlike traditional approaches that shuffle trial data, this method operates at the level of regression coefficients. For each session, a regression model is fitted to estimate session-specific beta coefficients, which capture the relationship between brain activity (here, state time courses) and the experimental condition. To test for significant changes across sessions, the method generates a (null) distribution by randomly permuting these beta coefficients across sessions (rather than permuting the data). This approach accounts for differences in session length, variations in condition proportions, and the lack of direct alignment between trials across sessions (as described in Supplementary Note S1). Similar to the across-trials test, this can be performed at each time point to produce a time-resolved statistical analysis.
The across-state-visits test, also novel in this context, evaluates whether the state time courses (represented by the Viterbi path, a discretised version of the state activation probabilities $\gamma_{tk}$) are associated with concurrently recorded physiological or behavioural signals over time, such as pupil size, heart rate, or skin conductance. For instance, one state might correspond to an increase in pupil size, while another might correspond to a decrease. The Viterbi path thus serves as the contrast, enabling the comparison of differences in the second set of signals. Standard permutation methods are however not suitable for this test, as shuffling time points would disrupt the temporal structure of the data. Instead, using a Monte Carlo approach, we generate surrogate Viterbi paths that preserve the original transition timings but randomly reassign which states are visited at each transition in a structured manner. This is done such that it maintains the statistical properties of the original data while breaking the observed association between states and the external signal, with further elaboration on this test available in Supplementary Note S2.
@@
Load and prepare data: Step 4 - Set-up and train an HMM (optional)
@@
Once the model is trained, the estimated parameters can be inspected to understand what each state represents and how the model behaves over time. These include the initial state probabilities (which reflect the state probabilities at the start of every segment of data), the average activation patterns (state-specific means), the covariance matrices (state-specific or state-averaged functional connectivity), and the transition probabilities (of transitioning from one state to another). The toolbox includes visualisation tools for these elements, making it easier to evaluate the fitted model before proceeding to statistical analysis.
@@
In addition to generating summary statistics for analysis, the fitted HMM can also be inspected directly. Users can access the estimated state means, covariance matrices, and transition probabilities to examine the spatial and temporal properties of each state. These outputs support model interpretation and quality control before proceeding to statistical testing.
@@
As stated, this protocol supports the following types of tests: across-subject, across-trials, across-sessions-within-subject, and across-state-visits.
@@
For across-subjects, across-sessions-within-subject, and across-trials-within-session tests, the protocol provides multivariate regression tests and univariate tests as well as Canonical Correlation Analysis (CCA). See Part 2 in Figure 2.
@@
For the Across-subjects, Across-sessions-within-subject, and Across-trials-within-session tests, the protocol includes the Non-Parametric Combination (NPC) algorithm to combine multiple p-values into fewer p-values with increased statistical power [24, 25]. Specifically, instead of getting a p-value for each pair of variables—that is a ($p\times q$) matrix of p-values—the NPC algorithm condenses these into either one p-value per row ($1\times p$ p-values), or one p-value per column ($1\times q$ p-values), or a single p-value for the entire test. In our implementation, we use Fisher’s method as the combining function, which efficiently aggregates the p-values while maintaining sensitivity to small values. This approach differs from CCA, which also produces a single p-value but does so by testing the strength of a multivariate relationship between two variable sets. In addition to statistical inference, CCA provides a latent representation of the data in the form of canonical variables—linear combinations that maximise correlation between the sets. NPC, by contrast, aggregates test results across multiple comparisons. While it offers flexibility and interpretability, it does not yield a latent representation or model the joint multivariate structure directly.
@@
The present protocol is designed to accommodate both task-based experimental designs and resting-state experiments. While many existing toolboxes such as such as FSL[28], SPM[29], AFNI[30], MNE[31], and CONN[32] provide robust support for standard group-level analyses, they are often limited to time-averaged representations of brain activity. Our protocol addresses this limitation by providing statistical inference on time-varying features (e.g., brain dynamics) and their relationship to behavioural or physiological variables.
What we refer to as across-subjects and across-trials analyses are supported by all the other toolboxes, enabling group-level inference and condition-based contrasts. However, support for longitudinal analyses in the way we present here (through the so-called across-sessions-within-subject testing) is lacking. Some tools (e.g., AFNI, CONN) allow users to combine multiple sessions by summarising each session separately (e.g., by computing average connectivity per session) and then comparing those summary metrics using group-level statistics. In contrast, our protocol retains the full temporal structure across sessions and enables trial-by-trial or timepoint-by-timepoint inference. This makes it possible to analyse how brain dynamics change from session to session and to ascertain when these changes take place as opposed to just whether they do. Additionally, our protocol includes across-state-visits testing, which allows users to link moment-by-moment occurrences of brain states to concurrently recorded behavioural or physiological variables (e.g., pupil size or heart rate) — a type of analysis not supported by the other toolboxes.
All of the compared tools support univariate and multivariate testing. However, CCA, a method well suited for linking multivariate neural features with multivariate behavioural data, is not natively supported in FSL, SPM, MNE, AFNI, or CONN. In some cases (e.g., FSL, SPM, MNE), CCA can be added using external tools such as PALM. In contrast, CCA is directly implemented in our framework.
The toolbox also supports combining results from multiple related tests, such as different cognitive measures, into a single test. This makes it possible to test whether patterns in the brain are linked to a broader behavioural profile rather than looking at each variable in isolation. While PALM supports test combination and could be used alongside other packages, this typically requires a manual setup. In contrast, our protocol integrates test combination directly into its core workflow, making it more accessible and easier to apply. Taking inspiration from PALM, our framework also provides hierarchical permutation testing to account for family relationships between subjects.
All toolboxes provide standard procedures for multiple testing correction (e.g., FWER and FDR) and cluster-based statistics. In our framework, these procedures are built directly into the main analysis pipeline, making them easier to apply without additional configuration.
Table 2 summarises the core differences in statistical testing capabilities across these toolboxes.
Category Feature FSL SPM MNE AFNI CONN
Type of Tests across-subjects Yes Yes Yes Yes Yes
across-trials Yes Yes Yes Yes Yes
@@
No No No No No
across-state-visits No No No No No
Methods Supported Multivariate Yes Yes Yes Yes Yes
Univariate Yes Yes Yes Yes Yes
@@
No No
one-state-vs-the-rest No No No No No
one-state-vs-another-state No No No No No
Test combination
Combine tests across
rows/columns/full matrix
No (default);
Yes via PALM
No (default);
Yes via PALM
No (default);
Yes via PALM
No No
@@
No No
Multiple Testing Correction & Cluster Statistics FWER Yes Yes Yes Yes Yes
FDR Yes Yes Yes Yes Yes
Cluster-based statistics Yes Yes Yes Yes Yes
Table 2: Comparison of statistical testing features across toolboxes. Each row represents a specific feature or analysis type, with ’Yes’ indicating that the toolbox supports the functionality and ’No’ indicating that it does not support the functionality.
@@
This protocol outlines pipelines for a comprehensive set of statistical tests, applicable to a broad range of scientific questions in neuroscience. We demonstrate these tests using publicly available data, as summarised below. In Protocol 1 (across-subjects), we analyse resting-state brain activity from 1,001 Human Connectome Project (HCP) participants across four sessions to examine its relationship to 15 traits related to cognitive performance, Supplementary Table 1. In Procedure 2 (across-trials), we study MEG data from a single person who participated in 15 sessions. During each session, the person watched both animate and inanimate objects while their brain activity in the occipital lobe was recorded. This analysis assesses differences in the brain responses when the person looks at animate objects compared to inanimate ones. In Procedure 3 (across-sessions-within-subject), we use the same dataset as in Procedure 2, but this time focus on changes over multiple sessions. This analysis shows whether the person exhibits changes in stimulus processing over time (i.e. across sessions) due for example, to learning, or if their brain representations remain stable. In Procedure 4 (across-state-visits), we analyse MEG data from 10 subjects scanned at rest in a dark room. During the scans, pupil size and brain activity were measured concurrently. Nine subjects completed two sessions, and one completed a single session. This analysis explores how changes in brain states, like the default mode network, relate to variations in pupil size[34, 35]. All data (except HCP) needed to reproduce the results of these workflows are hosted on Zenodo (DOI: 10.5281/zenodo.15213970), while the code is available at GitHub.
@@
  • GLHMM Python package: Install the GLHMM Python package and its dependencies using pip. The package is available for download at https://github.com/vidaurre/glhmm. This protocol is based on the latest release of GLHMM (version 1.1.1, released on July, 2025). The GLHMM toolbox is available both as a Python package and as GUI. The GUI allows users to run analyses through a user-friendly interface. A video tutorial ($\sim$30 minutes) demonstrating how to set up and use the GUI is also available; the link is provided in the GitHub repository.
@@ 1. Install and set up Python environment Timing: 1–5 min
    conda create --name glhmm_env python=3.10
@@
    Once the environment is activated, install the GLHMM package via pip:
    pip install glhmm
    This command installs the package along with all required dependencies. No additional setup is needed.
    The Jupyter notebooks used in this protocol are available on the GLHMM Protocols GitHub repository. These notebooks contain detailed examples for each of the four procedures described in this paper. Required datasets can be downloaded from Zenodo. However, the notebooks are designed to automatically download the data if it is not already present in the expected directories.
@@
    Behavioural data for 15 cognitive traits related to fluid intelligence is stored in data_cognitive_traits_HCP.npy and loaded into R_data, referred to as matrix $R$. Look at Supplementary Table 1 for the full list of traits. Additional confounds, such as sex, age, and fMRI head motion, are included in confounds_HCP.npy.
@@
    After training, the learned state properties and transition dynamics, such as initial state probabilities, state means, covariances, and transition probabilities, can be inspected to evaluate the model. These visualisations are included in the accompanying notebook.
@@
Procedure 3: Across-sessions-within-subject testing
Steps 1–5 follow the same process as in Procedure 2, as this protocol uses the same data and preprocessing setup. For details on data loading, preparation, and HMM training, refer to Procedure 2. This protocol starts at Step 6 and focuses on the statistical analysis for the across-sessions-within-subject test.
@@
All data used in this protocol are freely accessible on Zenodo (DOI: 10.5281/zenodo.15213970). The repository also includes a link to the associated code on GitHub: https://github.com/Nick7900/glhmm_protocols.
@@
All code is available on GitHub: https://github.com/Nick7900/glhmm_protocols, provided as Python notebooks that can be run directly in the cloud using Google Colab, so there is no need to install Python or any packages locally. For reproducibility, the repository is also archived on Zenodo: https://zenodo.org/records/15213970 (DOI: 10.5281/zenodo.15213970). This setup supports versioning for future updates, including new code and tutorials